OpenStudy (anonymous):

Can someone please help? Find the answer to -2+x=-3+4y solve for x and y. Thanks!

OpenStudy (anonymous):

You can do substitution or elimination. Substitution: -2+x=-3+4y Let's solve for x first x=-1+4y Now, substitute the equation(1+4y) as x in the original equation -2+(1+4y)=-3+4y

OpenStudy (anonymous):

Ok. I still can't figure it out. (don't accuse me of not trying)

OpenStudy (anonymous):

It's okay :) Um so -2+ (1+4y)=-3+ 4y Get the equation out of paratheses 1+4y-2=-3+4y Combine like terms. 4y-1= 4y-3 And now, you have learned that this problem has no solution. (No wonder you didn't get it) :D

OpenStudy (anonymous):

Oh. Ok. Thanks!

OpenStudy (anonymous):

Yeah :) Welcome.